40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1966     Entry time: Tue Oct 31 22:56:29 2017
Author: Craig 
Type: DailyProgress 
Category: TempCtrl 
Subject: Beatnote Fluctuations 

We have to figure out why our beatnote is flying around if we are going to do science down here.  It's preventing us from taking a low actuator noise spectrum off of our PLL, and that is required for measuring our true noise of our cavity beatnote.

Tonight I have set up the Agilent to always be taking spectra and recording them with acromag1.  This will allow us to track the beatnote motion and peak strength overnight with a data point coming once every 45 seconds. I'll have to write a peakTracker.py script which takes in a spectrum.txt and finds the peak, should be easy.

I have also written a channelLogger.py directly on acromag1, which is logging the channels for slow control voltages for both cavities, the in-loop and out-of-loop vaccan temp sensors, and an environmental temp sensors overnight.

The point of all of this is to look for obvious correlations between beatnote motion and temperature controls.


One thought I had today was the power difference in the laser cavities.  If I recall correctly, the North cavity has about 2.5 times as much power as the South cavity.  Also, our slow voltage laser temperature controls change both the laser frequency and the laser power.  This means that any large swings in the slow voltage laser control will result in significantly different power resonanting in the cavities, which could result in significant differential temperature changes.  Our cavity finesse is like 10000, so this effect is multiplied.  Perhaps we should think about evening out the power in the cavities, AKA putting a ND filter in front of the North cav?

ELOG V3.1.3-